A plane flies 955 km due east,then turns to north and flies another 649 km. Determine the polar coordinates of the plane's finam position


1
Expert's answer
2021-10-19T15:07:49-0400

The magnitude of the displacement vector:


"D=\sqrt{955^2+649^2}=1155\text{ km}.\\\space\\
\theta=\tan^{-1}\frac{955}{649}=55.8°\text{ E of N}."